OpenStudy (anonymous):

can someone explain this to me? find the integral of dx/(1-sinx)

OpenStudy (anonymous):

i don't know what that is

OpenStudy (anonymous):

try this: multiply and divide this integral in both denominator and numertr with (1+sinx) then you will have denominator as = 1-(sinx)^2 which is = (cosx)^2 and your numert will be= 1+sinx

OpenStudy (aravindg):

^^

OpenStudy (aravindg):

Good work @ijoshi .Welcome to Openstudy both of you :)

OpenStudy (anonymous):

thank you :)